# Neural Network Reprogrammability: A Unified Theme on Model Reprogramming, Prompt Tuning, and Prompt Instruction

### v2

@@ 2.1 From Vulnerability to Adaptation
Repurposing Sensitivity for Adaptation. This inherent input sensitivity, conventionally framed as a weakness, can however be reinterpreted as the enabler of neural network reprogrammability. The very mechanisms that allow adversarial examples to drastically alter model predictions–namely, the ability to find specific input directions that maximally change the output–can be repurposed. Thus, a seemingly detrimental vulnerability can be transformed beneficially. Instead of perturbing inputs to make a model fail (e.g., via gradient ascent on the loss function), we can engineer input transformations to guide a pre-trained model to perform tasks it was not originally designed or trained for. Such transformations eventually form the foundation of neural network reprogrammability
@@
This viewpoint marks a paradigm shift in model adaptation. While traditional transfer learning methods, e.g., fine-tuning (Pan and Yang, 2009) and domain adaptation (Long et al., 2016, 2018), focus on modifying model weights, with reprogrammability, models can be instead adapted by manipulating the data fed to a fixed function. This implies that the emphasis shifts from updating the model to transforming the data. Through careful design of transformations, we can effectively “reprogram” a model’s input-output mapping, without touching its core learned representations.
@@ 2.2 Formulation of Reprogrammability
The central mechanism for adapting the pre-trained model $f$ is input manipulation $I_{\lambda,\tau,\ell}:{\mathcal{X}}^{\rm T}\times{\mathcal{C}}\to{\mathcal{M}}$, which transforms target input ${\mathcal{X}}^{\rm T}\in{\mathcal{X}}^{\rm T}$ before it, or its subsequent representations, are processed by parts of $f$. The manipulation occurs at one or multiple of the model’s interfaces =${\mathcal{M}}=\{{\mathcal{X}}^{\rm S},{\mathcal{E}},{\mathcal{H}}\}$, which can be its raw input space ${\mathcal{X}}^{\rm S}$, its embedding space ${\mathcal{E}}$, or its deeper hidden representation spaces ${\mathcal{H}}$ [^1] When manipulations occur at embedding or hidden layers, it is helpful to view these layers as processing units whose own inputs are being altered.. Moreover, the manipulation can potentially be guided by contextual information ${\mathcal{C}}$, which refers to task-specific contexts and constraints guiding how the pre-trained model should be repurposed. This transformation is characterized by
@@
  • Linear Alignment. This also uses a matrix transformation $O_{\omega}({\mathbf{y}}^{\rm S})=\omega{\mathbf{y}}^{\rm S}$, but here $\omega\in{\mathbb{R}}^{|{\mathcal{Y}}^{\rm T}|\times|{\mathcal{Y}}^{\rm S}|}$ is a real-valued matrix. The matrix $\omega$ can be obtained either by gradient-based backpropagation (Tsao et al., 2024), e.g., optimizing a linear probe trained on top of the frozen pre-trained model’s outputs (Tsao et al., 2024); or statistically estimated without backpropagation, for example, by estimating the optimal probabilistic mapping matrix from observed statistics of source predictions and target labels (Cai et al., 2024a) [^2] We note that (Cai et al., 2024a) also relies on statistics and does not rely on gradient-based optimization. Still, here we distinguish it from the “statistical alignment” category as $\omega$ is not restricted to binary values and can represent more precise linear transformations..
@@ 2.3 Revisit Existing Manifestations of Reprogrammability
This represents an early structured approach (Chen, 2024; Elsayed et al., 2019) to repurpose pre-trained models without changing their weights, redirects model processing pathways through task-specific input perturbations to achieve desired outputs on a new target task. MR instantiates Eq. (1) by applying manipulation at either the input (for continuous data) or embedding interfaces (for discrete data), formally =$\ell=\{{\mathcal{X}}^{\rm S},{\mathcal{E}}\}$, with optimizing $\lambda\in\Lambda_{\rm learnable}$ against task-specific objective functions.
@@ 3 A Taxonomy of Reprogrammability
Table 2: A comparative analysis of existing RCA literature through the lens of reprogrammability. Under interface ($\ell$), setup ${\mathcal{X}}^{\rm S}$, ${\mathcal{E}}$, and ${\mathcal{H}}$ denote manipulations that occur in input space, embedding space, and hidden space, respectively. The columns under configuration ($\lambda$): setup OP and FX means optimizable and fixed manipulations. The columns under operator ($\tau$): setup AD, CO, and PR refers to additive, concatenative, and parametric operators. The columns under mapping ($\omega$): setup SA, LA, RA, and ID represent statistical alignment, linear alignment, rule-based (i.e., structured) alignment, and identity mapping.
@@ 3.2 Manipulation Location: Where to Modify
While MR rarely operates directly in hidden space, PT methods increasingly leverage these deeper modifications. P-Tuning v2 (Liu et al., 2021a) inserted learnable components between transformer layers, showing that deeper placement of prompts leads to more effective reprogramming. Multi-layer prompting like UniPELT (He et al., 2022) and UniAdapter (Lu et al., 2023) combined manipulations at strategic network depths for enhanced adaptation. Zhang et al. (2024c) explored the equivalence between visual prompts applied to internal activations and layer normalization.
@@ 3.3 Manipulation Operator: How to Transform
These operators extend inputs or representations by appending, prepending, or strategically inserting new elements. In MR, concatenative operators appear in approaches expanding input dimensionality (Wu et al., 2024a; Zhang et al., 2024b), such as by padding learnable patterns around target images (Wu et al., 2024a; Zhang et al., 2024b) or pretending learnable tokens to text input embeddings (Hambardzumyan et al., 2021). Concatenation dominates PT, from regular token prepending (Li and Liang, 2021; Lester et al., 2021) to dynamic insertion (Qin and Eisner, 2021). P-Tuning v2 (Liu et al., 2021a) concatenated prompts at multiple layers. Structured Prompting (Hao et al., 2022) incorporated domain knowledge into concatenated prompts. Concatenative operators also form the foundation of ICL (Brown et al., 2020; Wei et al., 2022b) in PI, particularly in CoT prompting (Wei et al., 2022b), where reasoning steps are appended to demonstrations. Moreover, the sensitivity of ICL to the order of concatenated examples has been analyzed (Min et al., 2022; Zhang et al., 2023b).
@@
These operators apply complex operations like projections or conditional normalization to inputs or intermediate representations: $I_{\lambda,{\rm param},\ell}({\mathbf{x}}^{\rm T})=\lambda_{c}({\mathbf{x}}^{\rm T})$, where $\lambda(\cdot)$ denotes the transformation function. In MR, parametric operators are used in approaches that apply non-linear transformations of input features for cross-domain adaptation (Tsai et al., 2020). For PT, parametric operators are strongly associated with adapter-based approaches (Houlsby et al., 2019; Wang et al., 2021; He et al., 2022) Within PI, parametric operators appear in retrieval-augmented (RAG) approaches (Lewis et al., 2020; Borgeaud et al., 2022; Trivedi et al., 2022; Dong et al., 2023), which transform inputs by incorporating retrieved documents.
@@ 3.4 Output Alignment Requirements
The simplest form of alignment occurs when model outputs directly correspond to target task outputs without additional processing. This applies when source and target output spaces are naturally compatible. Identity mapping is common in PI and PT for unimodal generative tasks, where the model’s output text directly serves as the task output. As models increase in capability, identity mapping becomes viable for increasingly complex tasks, where visual instruction following through direct generation is demonstrated without post-processing (Liu et al., 2023a, 2021b).
@@ 4.2 Implications of Taxonomic Position
We also summarize a few findings regarding the generalization of reprogrammed models. (1) Sample-specific Adaptation. Cai et al. (2024b) observed that manipulations conditioned on input characteristics $I_{\lambda,\tau,\ell}({\mathbf{x}}^{\rm T},c({\mathbf{x}}^{\rm T}))$ tend to generalize more effectively than static manipulations. This suggests dynamic adaptation better captures the underlying task structure rather than memorizing specific data patterns. (2) Representation Hierarchy Effects. The manipulation location influences adaptation capabilities. As layer depth increases, manipulations operate on more abstract representations (i.e., deeper layers) that generalize better across similar tasks but lose modality-specific details. This explains why manipulations targeting deeper network layers (particularly within transformer architectures (Liu et al., 2021a; He et al., 2022)) often achieve strong cross-domain transfer. (3) Scale-dependent Transfer. The generalization gap between input-level reprogramming and fine-tuning follows a power law relationship with model size (Bahng et al., 2022). That is, input-level reprogramming can underperform fine-tuning with smaller models but achieves competitive results as model size increases For models below a certain size threshold, fine-tuning consistently outperforms reprogramming. Above this threshold, reprogramming approaches competitive or superior performance with significantly fewer parameters (Wu et al., 2024a), e.g., from ResNet (He et al., 2016) to CLIP (Radford et al., 2021). Structured task decomposition improves generalization performance on new task-domain combinations (Andreas, 2022). Recently, Geng and Chen (2024) show that reprogramming is more robust than conventional fine-tuning with out-of-distribution data.
@@
Interface $\text{Interface Sensitivity}(f,\lambda)\approx\begin{cases}\epsilon&\text{if }|f|<S_{\text{threshold}}\\ \beta\cdot\log(|f|/S_{\text{threshold}})&\text{otherwise, }\end{cases}$
@@ 5.3 Limitations of Existing Programs
  • Data Efficiency: Although RCA avoids fine-tuning the entire model, learning effective “soft” prompts can still demand a considerable amount of target-task data, where the optimization of program parameters can be a challenging non-convex problem, requiring careful initialization and sufficient examples to converge to a good solution.
  • Task Complexity and Scope: Another limitation is that current RCA methods excel at tasks that align well with the pre-trained model’s inherent capabilities, such as classification, stylistic generation, or question answering, where the underlying knowledge is already present. However, for tasks requiring fundamentally new reasoning chains, complex symbolic manipulation not seen during pre-training (e.g., advanced multi-step mathematics for an LLM not extensively trained on it), or the acquisition of entirely novel skills, merely prepending a prompt or transforming inputs may be insufficient. In this case, expanding the model might be necessary.
  • Program Transferability: Moreover, a practical hurdle is that programs, especially learned soft prompts, are often highly specific to the particular model (and sometimes even the specific checkpoint) they were tuned for. A prompt optimized for one LLM architecture rarely transfers effectively to another, even one of similar size or family. This “hypersensitivity” restricts the reusability of prompts and forces re-optimization when models are updated or changed. While research into universal prompts or techniques for efficient prompt adaptation is ongoing, it remains a challenging frontier. Fixed, hard-coded prompts, though potentially more transferable, can be brittle and overly sensitive to minor variations in phrasing.
  • Optimization Challenges: MR and soft PT typically rely on gradient access to the model to optimize the prompt parameters. In many real-world scenarios where models are accessed via restricted APIs (black-box access), gradient-based optimization is impossible. This leaves downstream users with manual prompt engineering or less efficient derivative-free optimization methods for discovering effective prompts. Developing more powerful and sample-efficient black-box optimization techniques for prompt discovery in high-dimensional spaces is an open problem.
@@ 5.4 Ethical and Security Concerns
  • Evasion of Safeguards (i.e., prompt injection & jailbreaking): If anyone can repurpose a model via prompts, models could be coaxed into behaviors their creators did not intend (Chen et al., 2025). For example, LLMs might be reprogrammed via a cleverly crafted prompt to produce disallowed content, e.g., hate speech, even if safeguards are in place–a form of prompt injection attack (Liu et al., 2024c; Zou et al., 2023a). Ensuring that models cannot be easily reprogrammed for harmful purposes is an unsolved challenge.

### v3

@@ 2.1 From Vulnerability to Adaptation
Repurposing Sensitivity for Adaptation. This inherent input sensitivity, conventionally framed as a weakness, can however be reinterpreted as the enabler of neural network reprogrammability. The very mechanisms that allow adversarial examples to drastically alter model predictions–namely, the ability to find specific input directions that maximally change the output–can be repurposed. Thus, a seemingly detrimental vulnerability can be transformed beneficially. Instead of perturbing inputs to make a model fail (e.g., via gradient ascent on the loss function), we can engineer input transformations to guide a pre-trained model to perform tasks it was not originally designed or trained for. Such transformations eventually form the foundation of neural network reprogrammability.
@@
This viewpoint marks a paradigm shift in model adaptation. While traditional transfer learning methods, e.g., fine-tuning (Pan and Yang, 2009; Guo et al., 2022) and domain adaptation (Long et al., 2016, 2018), which focus on modifying weights of the source model, have been the de-facto practices, these approaches, especially the latter ones, find challenging source-free domain adaptation setting (Liang et al., 2020; Chi et al., 2021), where adaptation to a new target domain must be performed using only the pre-trained source model and target domain data, without any access to the original source data [^1] Many source-free methods still rely on updating parts of the model, such as batch normalization statistics or select layers, often through techniques like pseudo-labeling or entropy minimization (Dong et al., 2023b). This places them in a distinct category from the reprogrammability-based methods we discussed here.. By contrast, with reprogrammability, models can be instead adapted by manipulating the data fed to a fixed function. This implies that the emphasis shifts from updating the model to transforming the data. Through careful design of transformations, we can effectively “reprogram” a model’s input-output mapping, without touching its core learned representations.
@@ 2.2 Formulation of Reprogrammability
The central mechanism for adapting the pre-trained model $f$ is input manipulation $I_{\lambda,\tau,\ell}:{\mathcal{X}}^{\rm T}\times{\mathcal{C}}\to{\mathcal{M}}$, which transforms target input ${\mathcal{X}}^{\rm T}\in{\mathcal{X}}^{\rm T}$ before it, or its subsequent representations, are processed by parts of $f$. The manipulation occurs at one or multiple of the model’s interfaces ${\mathcal{M}}=\{{\mathcal{X}}^{\rm S},{\mathcal{E}},{\mathcal{H}}\}$, which can be its raw input space ${\mathcal{X}}^{\rm S}$, its embedding space $\mathcal{E}$}, or its deeper hidden representation spaces $\mathcal{H}$} [^2] When manipulations occur at embedding or hidden layers, it is helpful to view these layers as processing units whose own inputs are being altered.. Moreover, the manipulation can potentially be guided by contextual information $\mathcal{C}$}, which refers to task-specific contexts and constraints guiding how the pre-trained model should be repurposed. This transformation is characterized by
@@
  • Linear Alignment. This also uses a matrix transformation $O_{\omega}({\mathbf{y}}^{\rm S})=\omega{\mathbf{y}}^{\rm S}$, but here $\omega\in{\mathbb{R}}^{|{\mathcal{Y}}^{\rm T}|\times|{\mathcal{Y}}^{\rm S}|}$ is a real-valued matrix. The matrix $\omega$ can be obtained either by gradient-based backpropagation (Tsao et al., 2024), e.g., optimizing a linear probe trained on top of the frozen pre-trained model’s outputs; or statistically estimated without backpropagation, for example, by estimating the optimal probabilistic mapping matrix from observed statistics of source predictions and target labels (Cai et al., 2024a) [^3] We note that Cai et al. (2024a) also relies on statistics and does not rely on gradient-based optimization. Still, here we distinguish it from the “statistical alignment” category as $\omega$ is not restricted to binary values and can represent more precise linear transformations..
@@ 2.3 Revisit Existing Manifestations of Reprogrammability
This represents an early structured approach (Chen, 2024; Elsayed et al., 2019) to repurpose pre-trained models without changing their weights, redirects model processing pathways through task-specific input perturbations to achieve desired outputs on a new target task. MR instantiates Eq. (1) by applying manipulation at either the input (for continuous data) or embedding interfaces (for discrete data), formally $\ell=\{{\mathcal{X}}^{\rm S},{\mathcal{E}}\}$, with optimizing $\lambda\in\Lambda_{\rm learnable}$ against task-specific objective functions.
@@ 3 A Taxonomy of Reprogrammability
Table 2: A comparative analysis of existing RCA literature through the lens of reprogrammability. Under interface ($\ell$), setup ${\mathcal{X}}^{\rm S}$, $\mathcal{E}$}, and $\mathcal{H}$} denote manipulations that occur in input space, embedding space, and hidden space, respectively. The columns under configuration ($\lambda$): setup OP and FX means optimizable and fixed manipulations. The columns under operator ($\tau$): setup AD, CO, and PR refer to additive, concatenative, and parametric operators. The columns under mapping ($\omega$): setup SA, LA, RA, and ID represent statistical alignment, linear alignment, rule-based (i.e., structured) alignment, and identity mapping.
@@ 3.2 Manipulation Location: Where to Modify
While MR rarely operates directly in hidden space, PT methods increasingly leverage these deeper modifications. P-Tuning v2 (Liu et al., 2021b) inserted learnable components between transformer layers, showing that deeper placement of prompts leads to more effective reprogramming. Multi-layer prompting like UniPELT (He et al., 2022) and UniAdapter (Lu et al., 2023) combined manipulations at strategic network depths for enhanced adaptation. Zhang et al. (2024c) explored the equivalence between visual prompts applied to internal activations and layer normalization.
@@ 3.3 Manipulation Operator: How to Transform
These operators extend inputs or representations by appending, prepending, or strategically inserting new elements. In MR, concatenative operators appear in approaches expanding input dimensionality (Wu et al., 2024a; Zhang et al., 2024b), such as by padding learnable patterns around target images (Wu et al., 2024a; Zhang et al., 2024b) or pretending learnable tokens to text input embeddings (Hambardzumyan et al., 2021). Concatenation dominates PT, from regular token prepending (Li and Liang, 2021; Lester et al., 2021) to dynamic insertion (Qin and Eisner, 2021). P-Tuning v2 (Liu et al., 2021b) concatenated prompts at multiple layers. Structured Prompting (Hao et al., 2022) incorporated domain knowledge into concatenated prompts. Concatenative operators also form the foundation of ICL (Brown et al., 2020; Wei et al., 2022b) in PI, particularly in CoT prompting (Wei et al., 2022b), where reasoning steps are appended to demonstrations. Moreover, the sensitivity of ICL to the order of concatenated examples has been analyzed (Min et al., 2022; Zhang et al., 2023b).
@@
These operators apply complex operations like projections or conditional normalization to inputs or intermediate representations: $I_{\lambda,{\rm param},\ell}({\mathbf{x}}^{\rm T})=\lambda_{c}({\mathbf{x}}^{\rm T})$, where $\lambda(\cdot)$ denotes the transformation function. In MR, parametric operators are used in approaches that apply non-linear transformations of input features for cross-domain adaptation (Tsai et al., 2020). For PT, parametric operators are strongly associated with adapter-based approaches (Houlsby et al., 2019; Wang et al., 2021; He et al., 2022) Within PI, parametric operators appear in retrieval-augmented (RAG) approaches (Lewis et al., 2020; Borgeaud et al., 2022; Trivedi et al., 2022; Dong et al., 2023a), which transform inputs by incorporating retrieved documents.
@@ 3.4 Output Alignment Requirements
The simplest form of alignment occurs when model outputs directly correspond to target task outputs without additional processing. This applies when source and target output spaces are naturally compatible. Identity mapping is common in PI and PT for unimodal generative tasks, where the model’s output text directly serves as the task output. As models increase in capability, identity mapping becomes viable for increasingly complex tasks, where visual instruction following through direct generation is demonstrated without post-processing (Liu et al., 2023a, 2021c).
@@ 4.2 Implications of Taxonomic Position
We also summarize a few findings regarding the generalization of reprogrammed models. (1) Sample-specific Adaptation. Cai et al. (2024b) observed that manipulations conditioned on input characteristics $I_{\lambda,\tau,\ell}({\mathbf{x}}^{\rm T},c({\mathbf{x}}^{\rm T}))$ tend to generalize more effectively than static manipulations. This suggests dynamic adaptation better captures the underlying task structure rather than memorizing specific data patterns. (2) Representation Hierarchy Effects. The manipulation location influences adaptation capabilities. As layer depth increases, manipulations operate on more abstract representations (i.e., deeper layers) that generalize better across similar tasks but lose modality-specific details. This explains why manipulations targeting deeper network layers (particularly within transformer architectures (Liu et al., 2021b; He et al., 2022)) often achieve strong cross-domain transfer. (3) Scale-dependent Transfer. The generalization gap between input-level reprogramming and fine-tuning follows a power law relationship with model size (Bahng et al., 2022). That is, input-level reprogramming can underperform fine-tuning with smaller models but achieves competitive results as model size increases For models below a certain size threshold, fine-tuning consistently outperforms reprogramming. Above this threshold, reprogramming approaches competitive or superior performance with significantly fewer parameters (Wu et al., 2024a), e.g., from ResNet (He et al., 2016) to CLIP (Radford et al., 2021). Structured task decomposition improves generalization performance on new task-domain combinations (Andreas, 2022). Recently, Geng and Chen (2024) showed that reprogramming is more robust than conventional fine-tuning with out-of-distribution data (Jiang et al., 2023; Zheng et al., 2023; Han et al., 2025).
@@
[…]
@@ 5.3 Limitations of Existing Programs
  • Data Efficiency: Although RCA avoids fine-tuning the entire model, learning effective “soft” prompts can still demand a considerable amount of target-task data, where the optimization of program parameters can be a challenging non-convex problem, requiring careful initialization and sufficient examples to converge to a good solution. Future works may draw inspiration from meta-learning methods (Finn et al., 2017; Garnelo et al., 2018a, b; Ye and Yao, 2022; Du et al., 2023; Ye et al., 2023b, a; Liu et al., 2021a) and domain generalization (Zhong et al., 2024), to reduce required sample complexity.
  • Task Complexity and Scope: Another limitation is that current RCA methods excel at tasks that align well with the pre-trained model’s inherent capabilities, such as classification, stylistic generation, or question answering, where the underlying knowledge is already present. However, for tasks requiring fundamentally new reasoning chains, complex symbolic manipulation not seen during pre-training (e.g., advanced multi-step mathematics for an LLM not extensively trained on it), or the acquisition of entirely novel skills, merely prepending a prompt or transforming inputs may be insufficient. In this case, it is worthwhile to incorporate complex reasoning strategies, e.g., causal reasoning (Chi et al., 2024), into the reprogramming pipeline.
  • Program Transferability: Moreover, a practical hurdle is that programs, especially learned soft prompts, are often highly specific to the particular model (and sometimes even the specific checkpoint) they were tuned for. A prompt optimized for one LLM architecture rarely transfers effectively to another, even one of similar size or family. When the source model is needed to be adapted for multiple distributionally similar target tasks, this “hypersensitivity” restricts the reusability of prompts and forces re-optimization for each new task. While research into universal prompts or techniques for efficient prompt adaptation is ongoing, it remains a challenging frontier. Hard-coded prompts, though potentially more transferable, can be brittle and overly sensitive to minor variations in phrasing.
  • Optimization Challenges: MR and soft PT typically rely on gradient access to the model to optimize the prompt parameters. In many real-world scenarios where models are accessed via restricted APIs (black-box access), gradient-based optimization is impossible. This leaves downstream users with manual prompt engineering or less efficient derivative-free optimization methods, such as evolutionary algorithms-based ones (Jiang et al., 2016), for discovering effective prompts. Developing more powerful and sample-efficient black-box optimization techniques for prompt discovery in high-dimensional spaces is an open problem.
@@ 5.4 Ethical and Security Concerns
  • Evasion of Safeguards (i.e., prompt injection & jailbreaking): If anyone can repurpose a model via prompts, models could be coaxed into behaviors their creators did not intend (Chen et al., 2025). For example, LLMs might be reprogrammed via a cleverly crafted prompt to produce disallowed content, e.g., hate speech, even if safeguards are in place–a form of prompt injection attack (Liu et al., 2024c; Zou et al., 2023a). On the other hand, the reprogrammed models do not possess built-in robustness against malicious manipulations, e.g., adversarial and backdoor attacks (Chen et al., 2025). Ensuring that models cannot be easily reprogrammed for harmful purposes via building bespoke defenders, e.g., adversarial detectors (Zhang et al., 2025) and purifiers (Sun et al., 2025), is an important yet unresolved challenge.